You sell a put option on one share of stock. The put has a premium of $4 and a strike/exercise price of $98. The stock currently has a price of $101.20 per share. On the day that the option expires, the stock is selling for $94. What ends up being your net payoff on this position?
Answer #1

Loss from put option expiring in the money = final price - strike price = 94 - 98 = $4

option premium earned = $4

Hence total gains made = option premium + loss from put option = 4 - 4 = $0